# Fisher information, Crámer-Rao complexity and temperature in logistic map

### v1

@@ III.3 Temperature associated to logistic map
where $\{x_{k}^{(j)}\}$ is the orbit starting at $x_{0}^{(j)}$ with $j=1,\ldots,M$.
III.3.1 Numerical results
We calculate MP (16) for the same range of values of $\mu=0.05,0.1,\ldots,3.95,4$ as before, with $M=10^{5}$ initial conditions and $N=10^{3}$. It is worth to mention that for $M>10^{5}$ the MP remained with the same behavior. When calculating MP from definition (16), we observed the existence of a short time transient along with the presence of small variations of the $T(\mu,N,M)$ for some values of $\mu$. In order to define a temperature for a large number of steps we have taken an average neglecting the transient. The dependence of the MP in function of $\mu$ is shown in Fig. 4. For a better visualization, the MP was normalized according to $T(\mu,N,M)/\max\{T(\mu,N,M)\}$ with the the maximum taken for $N\geq N_{0}$, being $N\leq N_{0}$ the transient region.
Figure 4: Averaged value of the map temperature (16) for $\mu=0.05,0.1,\ldots,3.9,3.95,4$ with $N=10^{3}$ number of steps and $M=10^{5}$ initial conditions.
III.3.2 Temperature as a function of time
III.3.3 Some formal results
We can demonstrate the following formal result for the map temperature MP (16).
Theorem: For the function $T(\mu,N)$ we have:
  • $(I)$
    If $0<\mu\leq 3$ then $T(\mu,N)\rightarrow 0$ when $N\rightarrow\infty$.
  • $(II)$
    If $\mu=4$ then $|T(4,N)-K|<\varepsilon$ for all $N\geq C$ with $C>0$.
Proof: In order to bound $\cos(2^{n+1}\arccos{(1-x_{0}^{(j)})})-\cos(2^{n}\arccos{(1-x_{0}^{(j)})})]^{2}$ we note that this expression can be recasted as $(\cos(2(2^{n}y_{j}))-\cos(2^{n}y_{j}))^{2}$ with $y_{j}=\arccos{(1-x_{0}^{(j)})}$. In turn, we have $(\cos(2(2^{n}y_{j}))-\cos(2^{n}y_{j}))^{2}=(1/4)(e^{i(2(2^{n}y_{j}))}-e^{-i(2(2^{n}y_{j}))}-(e^{i(2^{n}y_{j})}-e^{i(2^{n}y_{j})}))^{2}$
@@ III.4 Fisher information and CR complexity as a function of temperature
In […] it has been shown that the Fisher information of the Liouville distribution $\rho(q,p)$ of the canonical ensemble is a decreasing monotonic function of temperature. Inspired by this feature, now we study the FI and the CR as functions of the temperature. In Fig.
III.5 Fisher information and CR complexity as a function of time
@@ IV Conclusions
The results of this work are potentially extended to other chaotic maps, for instance the tent map and the circular map.
@@ Acknowledgments
Samuel S. Santos, Guilherme V. B. Junior and Ignacio S. Gomez acknowledge support from the Department of Exact and Natural Sciences of the State University of Southwest Bahia (UESB), Itapetinga, Bahia, Brazil. Ignacio S. Gomez acknowledges support received from the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grant Number 316131/2023-7. Guilherme Vieira Brito Junior acknowledges support received from FAPESB as a scientific initiation scholarship.

### v2

@@ III.3 A temperature associated to the logistic map
where $\{x_{k}^{(j)}\}$ is the orbit starting at $x_{0}^{(j)}$ with $j=1,\ldots,M$. The physical meaning of the expression (18) is that $T(\mu,N,M)$ characterizes the transition dynamics of the logistic map at the $N$-th step, averaged over the $M$ initial conditions belonging to the space state $[0,1]$.
We calculate MP (18) for the same range of values of $\mu=0.05,0.1,\ldots,3.95,4$ as performed in Section III.A, with $M=10^{5}$ initial conditions and $N=10^{3}$. It is worth to mention that for $M>10^{5}$ the MP remained with the same behavior. When calculating MP from definition (18), we observed the existence of a short time transient along with the presence of small variations of the $T(\mu,N,M)$ for some values of $\mu$. In order to define a temperature for a large number of steps we have taken the average of (18) neglecting the transient. The dependence of the MP in function of $\mu$ is shown in Fig. 6. For a better visualization, the MP was normalized according to $T(\mu,N,M)/\max\{T(\mu,N,M)\}$ with the the maximum taken for $N\geq N_{0}$, being $N\leq N_{0}$ the transient region.
Figure 6: The map temperature (18) for the values of the map parameter $\mu=0.05,0.1,\ldots,3.9,3.95,4$ with $N=10^{3}$ number of steps and $M=10^{5}$ initial conditions is illustrated.
We also can analyze the behavior of the map temperature $T(\mu,N,M)$ as a function of the time step, which is illustrated in Fig. 7. It is observed that an associated temperature is well defined for the regimes with less complexity, that is, the regular $\mu=1$ and the fully chaotic $\mu=4$ ones. By contrast, when the regime presents more complexity, $\mu=3.56695$ in our case, the temperature exhibits non trivial fluctuations that prevents to define it. In turn, this can be understood from the manifestation of the complex behavior of the onset of chaos.
Figure 7: Map temperature $T(\mu,N,M)$ versus time step of the logistic map for the parameter values $\mu=1$ (upper), $\mu=3.56695$ (center) and $\mu=4$ (bottom) provided with $M=10^{5}$ initial conditions.
@@ III.4 Fisher information and CR complexity as a function of temperature
Using the Liouville canonical distribution $\rho(q,p)$ expressed as a product of the position and momentum distributions, in [14] it has been shown that the Fisher information of $\rho(q,p)$ is a decreasing monotonic function of temperature. Inspired by this study and considering that the MP reaches an equilibrium value as if they were $M$ particles in contact with a thermal reservoir, now we investigate the FI and the CR complexity as functions of the temperature. From the upper panel of Fig. 8 we see that the Fisher information behaves predominantly within a concentrated strain of values when the temperature is near to the maximum $T=1$. We observe a peak of the FI near to the zero temperature, corresponding to the regular dynamics. Also, greater values of the FI are matched with more regular regimes. All this seems to be in agreement with the fact that nonzero map temperatures must be associated to a dynamics with some level of chaos, in order to be consistent with the idea of individual trajectories wandering over the all state space, thus corresponding to lower values of the Fisher information. Contrarily, for regular regimes $0<\mu\leq 3$ the orbits converge to a limit $x_{\infty}$, thus implying a null average velocity $x_{N+1}^{(j)}-x_{N}^{(j)}\sim 0$ in the limit $N\rightarrow\infty$ of large iterations for all initial condition $j=1,\ldots,M$ and then resulting in a zero temperature $T(\mu,N,M)$ and in a greater value of the Fisher information. From the lower panel in Fig. 8, we see consistently that lower values of the CR complexity are associated to the regimes with a simple dynamic like the regular and fully chaotic ones. The peak in the region $0.8\leq T\leq 1$ corresponds to the onset of chaos zone, thus expressing the effectiveness of the CR complexity in detecting complex behavior.
Figure 8: Fisher information and Crámer-Rao complexity versus normalized average temperature (18) of the logistic map after $N=10^{3}$ time steps and for $M=10^{5}$ initial conditions.
@@ IV Conclusions
We have presented a study of the logistic map from FI and CR complexity numerical calculations for the invariant density distribution and, by employing an analogy with the Equipartition Theorem, we have defined an associated map temperature. Our contributions are enumerated as follows.
  • (a)
    The Fisher information and the Crámer-Rao complexity of the invariant density allow to distinguish regimes and transition dynamics in terms of the parameter of the logistic map, as evidenced in Figs. 1 and 2. In particular, the FI is useful for detecting regular regimes expressed by peaks due to its sensitivity, while the CR complexity characterizes regular, chaotic and complex regimes with a presence of peaks in transitional regions like the onset of chaos, Pameau-Maneville scenario. In addition, the graph FI versus CR complexity can be employed for detecting a correlation between them in order to identify regions of complex behavior (Fig. 3).
  • (b)
    The time evolution of the FI and of the CR complexity of the invariant density characterize several features of the dynamics for representative values of the parameter of the map, as we can see from Figs. 4 and 5, and provided with a small number of iterations. For the regular regime $\mu=1$, the presence of intermittencies with a decreasing and an abrupt grow represent a violation of the Frieden’s informational intepretation of the Second Law (17). The CR complexity for $\mu=1$ expresses a decreasing monotonic behavior, thus suggestively indicating that when the Frieden’s condition $dI/dt\leq 0$ is dissatisfied, the CR complexity still preserves the Frieden’s interpretation with $dC/dt\leq 0$ and in agreement with the fact that regular regimes must correspond to a vanishingly small complexity. When $\mu=3.56695$ and $\mu=4$, fluctuations and oscillations are manifested both for the FI and for the CR complexity, indicating that these statistical quantifiers do not fully stabilize during chaotic behavior.
  • (c)
    The map temperature MP proposed in (18), arising from an analogy between the velocities of the particles of a gas and the mean velocity between two consecutive positions for several initial conditions, is a good quantifier of the dynamics, as we can see from Fig. 6. With low number of iterations $N=10^{3}$, we observe that for all the regular regime $0<\mu\leq 3$ the map temperature is zero, consistent with the convergent orbits for all initial condition. In the region $3<\mu\leq 4$ the MP is nonzero with a non monotonic from $\mu=3$ until $\mu=4$ with the presence of variations in the complex regime (Pameau-Maneville scenario) $3.56995\leq\mu\leq 3.82843$ and then decreasing up to $\mu=4$. The MP in function of the time has evidenced an asymptotic stabilization only for regimes with low complexity, i.e. for $\mu=1$ (regular) and $\mu=4$ (fully chaotic) while for the onset of chaos $\mu=3.56695$ the fluctuations impede the reaching of an equilibrium temperature.
  • (d)
    From the MP (18), we have analyzed the FI and the CR complexity as functions of the temperature in light of [14], illustrated in Fig. 8. By means of an analogy between particles in the canonical ensemble at an equilibrium temperature and the MP originated by the dynamics of the initial conditions, we have found that for the invariant density of the logistic map the condition $dI/dT\leq 0$ is not guaranteed, mainly due the complex dynamics of the transition region (onset of chaos) corresponding to the agglomerated points in $0.8\leq T\leq 1$. Maximum values of the FI and the CR complexity detect temperatures where the behavior is regular and complex.
The present study of the logistic map by means of the FI and CR complexity of the invariant density, along with our definition of map temperature, has been shown to be consistent in characterizing regular, fully chaotic and complex regimes, with an additional advantage of requiring smaller number of iterations and of initial conditions than other ones employed in the literature. A second benefit of our study is the theory of Frobenius Perron and Koopman operators of the densities connecting the numerical calculations of the invariant density. Finally, a third advantage is that the MP only requires the calculation of the mean kinetic energy for a relatively small number of iterations in order to display dynamical features. The main results of our study are illustrated in a summarized way in Table 1. In future researches we hope to apply the formalism presented in other chaotic maps [22, 23, 24].
parameter $\mu$ $T(\mu,N,M)$ time $t$
Fisher localized delimited Frieden’s
information at $\mu=1$ in a strip condition (17)
$I[p]$ (11) and asymptotically
concentrated with
around the $dI/dt\sim 0$ and
maximum fluctuations,
value $T=1$ but violated
for $\mu=1$
Crámer-Rao sensitive neither asymptotically
complexity variations decreasing constant
$C[p]$ (12) in $3\leq\mu\leq 4$ nor with
increasing $dC/dt\sim 0$
map $\sim 0$ until zero for
temperature $\mu=3$, where $\mu=1$ and
$T(\mu,N,M)$ it starts to nonzero for
(18) grow, variations $\mu=4$ with
in $3\leq\mu\leq 4$ fluctuations
for $\mu=3.56695$
Table 1: A summary of the main characteristics of the Fisher information, CR complexity and map temperature of the logistic map as functions of the parameter $\mu$, the map temperature $T(\mu,N,M)$ and the time step $t$.
@@ Acknowledgments
Samuel S. Santos, Guilherme V. B. Junior, Ignacio S. Gomez and Ronaldo Thibes acknowledge support from the Department of Exact and Natural Sciences of the State University of Southwest Bahia (UESB), Itapetinga, Bahia, Brazil. Ignacio S. Gomez acknowledges support received from the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grant Number 316131/2023-7. Guilherme Vieira Brito Junior acknowledges support received from FAPESB as a scientific initiation scholarship.